### v2

# High-resolution Giant Metrewave Radio Telescope Hi 21 cm imaging of the host galaxy of FRB 20250316A

Abstract
We report Giant Metrewave Radio Telescope (GMRT) Hi 21cm imaging of NGC 4141, the host galaxy of FRB 20250316A at . Our GMRT Hi 21cm images have spatial resolutions, at , of kpc, and find evidence for (i) a companion galaxy, LEDA 2582852, to the south-west, (ii) a nearby (27-kpc distant) Hi cloud to the south-west, (iii) disturbances in the Hi distributions of both NGC 4141 and LEDA 2582852, and (iv) high Hi column densities in the south-western outskirts of NGC 4141. A Sloan Digital Sky Survey spectrum yields a low metallicity and a high star-formation rate (SFR) surface density in the south-western (SW) disk of NGC 4141, and an H-based SFR estimate that is significantly higher than that, at the same location, from the Galaxy Evolution Explorer near-ultraviolet image, indicating a recent burst of star-formation. The total SFR of NGC 4141 is also found to be higher via the H line than from the 1.4 GHz radio continuum. The above evidence indicates that NGC 4141 has recently (within the last Myr) acquired metal-poor gas, via either a merger or accretion, that resulted in the south-western starburst and that may have also triggered large-scale star-formation activity in NGC 4141, resulting in the formation of the stellar progenitor of FRB 20250316A and the other transients. Our highest-resolution (480 pc) GMRT Hi 21 cm image finds no Hi 21 cm emission from the location of FRB 20250316A or the nearby star-forming region, suggesting that most of the Hi here has been either ionized or converted into the molecular phase. Our non-detection of continuum emission at the location of FRB 20250316A yields the upper limit erg s-1, on the 1.38 GHz radio luminosity of a putative persistent radio source associated with FRB 20250316A, one of the strongest constraints on the radio luminosity of such an associated persistent radio source.
I Introduction
Fast Radio Bursts (FRBs) are millisecond-duration radio transient sources, with an isotropic sky distribution and high dispersion measures, indicating that they are extragalactic objects (e.g. Lorimer et al., 2007; CHIME/FRB Collaboration, 2021; Law et al., 2024a; Chime/FRB Collaboration, 2025a; FRB Collaboration, 2026; Shannon et al., 2025). More than a thousand FRBs are known today, with tens of sources showing repeated bursts with similar properties (e.g. Spitler et al., 2016; CHIME/FRB Collaboration, 2019, 2023). Radio interferometric studies have yielded accurate localizations and host galaxy identifications for FRBs out to (e.g. Chatterjee et al., 2017; Bannister et al., 2019; Ravi et al., 2019; Marcote et al., 2020; Ryder et al., 2023; Law et al., 2024a; Caleb et al., 2025). The majority of these FRBs with high-probability associations arise in star-forming galaxies (e.g. Heintz et al., 2020; Gordon et al., 2023a; Sharma et al., 2024), although a few have been detected in massive ellipticals and even a globular cluster (e.g. Kirsten et al., 2022; Sharma et al., 2023; Shah et al., 2025).
Despite the large number of known FRBs, their origin remains a mystery today. The discovery of an FRB-like burst from a Galactic magnetar (Bochenek et al., 2020; CHIME/FRB Collaboration, 2020) implicates highly-magnetized neutron stars as one channel for FRB progenitors (e.g. Metzger et al., 2019; Margalit et al., 2019). However, there remain a plethora of proposed emission mechanisms and sources which need ancillary constraints to reveal the predominant origins of FRBs. One such path is to study the local burst environment which describes the physical conditions that give rise to FRBs and therefore offers insight into their origin.
Unfortunately, the large distances of most FRBs make it difficult to examine their environments with high physical spatial resolution. Recently, Chime/FRB Collaboration (2025b) reported the detection of the brightest FRB ever, FRB 20250316A, with a very low dispersion measure, pc cm-3, indicating a nearby object. The CHIME/FRB Outrigger data yielded a localization with an accuracy of mas, coincident with NGC 4141, a star-forming galaxy at (Chime/FRB Collaboration, 2025b; Springob et al., 2005). The very low redshift of NGC 4141 and the accurate localization imply that the position of FRB 20250316A is known to an accuracy of pc,111Throughout this paper, we will use a Tully-Fisher distance of 33.2 Mpc for NGC 4141 (Tully et al., 2023). i.e. similar to the size of a molecular cloud. Here, we report high-resolution Giant Metrewave Radio Telescope (GMRT) Hi 21cm imaging of NGC 4141, to study the neutral atomic gas conditions that gave rise to FRB 20250316A.
II Observations and Data Analysis
| Angular Resolution | Spatial Resolution | RMS noise | NHI () |
| () | kpc kpc | (mJy beam-1) | (cm-2) |
| 0.65 | |||
| 0.49 | |||
| 0.45 | |||
| 0.38 | |||
| 0.30 | |||
| 0.25 |
We used the GMRT Band-5 receivers to image the Hi 21cm emission from NGC 4141 on 2025 March 31, with 7 hours of on-source time (proposal DDTC429; PI: N. Kanekar). The GMRT Wideband Backend was used as the correlator, with a bandwidth of 200 MHz subdivided into 8,192 channels and covering the frequency range MHz. This yielded a velocity resolution of km s-1 at the redshifted Hi 21cm line frequency of MHz. Observations of the compact source 1313+549, for 5m after every 40m on NGC 4141, were used to measure the antenna-based complex gains and bandpass shapes, while the flux density scale was determined via observations of 3C286.
The data were analysed in casa (version 6.6.5; CASA Team, 2022), following standard procedures. After initial data editing to remove non-working antennas and data affected by radio frequency interference (RFI), the antenna-based complex gains and bandpass shapes were measured using the routines gaincalR and bandpassR (Chowdhury et al., 2020, 2022). These gains were applied to the multi-channel target source visibilities, and the calibrated visibilities then split out. After further editing of data affected by RFI, using a combination of manual flagging and the package aoflagger (Offringa et al., 2012), the visibility data were averaged to a resolution of MHz and then imaged using the task tclean. The imaging used the w-projection and MT-MFS algorithms (Cornwell et al., 2008; Rau and Cornwell, 2011), to image a region of size , well beyond the first null of the GMRT Band-5 primary beam. Following the imaging, we used a standard self-calibration procedure to refine the antenna-based complex gains, using 4 rounds of phase-only self-calibration and imaging, followed by 1 round of amplitude-and-phase self-calibration and imaging. The imaging during the self-calibration procedure used Briggs weighting, with robust parameter of -0.5, to obtain a narrower synthesized beam (). We then subtracted out the continuum image from the calibrated visibilities, using the task uvsub, and carried out one more round of data editing with aoflagger, based on the residual visibilities, before repeating the amplitude-and-phase self-calibration and making the final continuum image. This image, made with Briggs weighting and robust parameter of +1 (for higher sensitivity), has an angular resolution of and an RMS noise of Jy/Beam near the image centre. No continuum emission is detected from the FRB location; the upper limit on the 1.38 GHz FRB continuum flux density is Jy.
We next applied the antenna-based gains to the spectral-line visibilities, and then subtracted out the final continuum image from the calibrated visibilities, again using the task uvsub. A final round of aoflagger was then run on the residual spectral-line visibilities, with lower thresholds to remove weak RFI. The continuum model was then added back to the residual visibilities, to produce the final calibrated spectral-line dataset.
The GMRT contains baselines of length km, and is thus sensitive to Hi 21cm emission from a wide range of angular scales. We produced spectral cubes at a range of angular resolutions, , to capture both the fine structure and the large-scale features. For each angular resolution, we initially made a continuum image at the same resolution and subtracted this out from the calibrated spectral-line visibilities, using the routine uvsub. We then Hanning-smoothed and resampled the residual visibilities, to obtain a spectral resolution of km s-1. Finally, for each resolution, we imaged the residual visibilities with appropriate tapering of the long baselines and Briggs weighting (with robust values of either 0.0 or +0.5) to produce the Hi 21cm cube. All spectral cubes were made at a velocity resolution of km s-1. The cubes were cleaned down to a threshold of , where is the per-channel RMS noise for each cube. The parameters of the different Hi 21cm spectral cubes are listed in Table 1.
We fitted a single-component Gaussian model to measure the 1.38 GHz flux density of NGC 4141 from the -resolution continuum image (to avoid resolving out the extended continuum). This yielded a 1.38 GHz flux density of mJy. Similar values were obtained from the continuum images at resolutions of and .
We compared the positions of unresolved sources in our highest-resolution continuum image (resolution ) with their positions in the Very Large Array FIRST 1.4 GHz image (Becker et al., 1995), using single-Gaussian models to identify the source position. The source positions in the two images were found to be consistent within . This is a conservative estimate of the astrometric accuracy of the GMRT images and spectral cubes.
Finally, we used the smooth and clip algorithm (Serra et al., 2012) in the source-finder software SoFiA (Serra et al., 2015; Westmeier et al., 2021) to identify Hi 21cm emission features in the spectral cubes and generate the Hi 21cm moment images. This algorithm compares the number of positive and negative features in each spectral cube at a range of spatial and spectral resolutions, to identify “reliable” positive emission features. For each cube, our search used Gaussian spatial kernels of 0, 3, and 6 pixels (with the synthesized beam of each cube sampled by 5 pixels), and boxcar spectral kernels of 0, 3, 5, 7, and 15 channels. We used a detection threshold of either or for the different cubes (chosen to obtain a sufficiently large number of negative features, to ensure accurate reliability estimates), and a high reliability threshold of 0.9, to ensure that any identified emission features are likely to be real. We note that this is a conservative approach, and that there could be weak features in the cubes that are not identified at this reliability threshold. Conversely, this has the advantage that any identified features are very likely to be real. We note that, for the resolution cube, we experimented with source-finding with SoFiA at lower reliability thresholds, down to a threshold of 0.7; no other candidate line emitters were detected in these searches within the FWHM of the GMRT primary beam. Finally, for all emission regions identified by the SoFiA-based searches, we extracted Hi 21cm spectra from the cubes to confirm the reality of the features (see Fig. 2).
III Results and Discussion
FRB 20250316A is located at R.A.=12h09m44.319s, Dec.=+58d50m56.708s, with an uncertainty of mas (Chime/FRB Collaboration, 2025b). This position lies within the disk of NGC 4141, an SBc-type spiral galaxy at . Besides FRB 20250316A, a number of transients have been discovered in this galaxy over the last two decades, including two Type-II supernovae, 2009E and 2008X (e.g. Salo et al., 2025), an optical transient AT 2025erx (Andreoni and Palmese, 2025), and an X-ray source recently localized with the Chandra X-ray Observatory (Sun et al., 2025a, b). The number of detections of transient sources in NGC 4141 suggests a recent increase in the star-formation activity in the galaxy.
We used the GMRT 300-resolution cube to carry out a wide-field search for Hi 21cm emission in and around NGC 4141, as this provides a good balance between Hi column density sensitivity and image fidelity (note that the 500-resolution image is made from a relatively small number of short GMRT baselines, which are more susceptible to systematic errors, e.g. due to RFI). The top panel of Figure 1 shows the velocity-integrated Hi 21cm moment image at an angular resolution of 300, obtained from our SoFiA smooth-and-clip search, with the coordinates relative to the location of NGC 4141. Three distinct emission regions are visible here; the three lower panels of the figure show a zoom-in on each region, with the Hi 21cm emission shown in contours overlaid on the DECaLS g-band image (Dey et al., 2019), in colourscale. Hi 21cm emission is detected from [A] the FRB host galaxy, NGC 4141, [B] an Hi cloud that is (i.e. kpc) to the south-west of NGC 4141, and [C] a companion galaxy, LEDA 2582852, (i.e. kpc) to the south-west of NGC 4141. The peak Hi column density of the south-western Hi cloud is cm-2; this cloud does not have an optical counterpart in the DECaLS g-band image. We note that all three emission regions are also clearly detected in the 500-resolution cube.
The three panels of Figure 2 show the Hi 21cm spectra from [A] NGC 4141, [B] the Hi cloud, and [C] LEDA 2582852, with Hi 21cm line flux density plotted against barycentric velocity, in km s-1, relative to . The Hi 21cm emission from both the Hi cloud and LEDA 2582852 is seen to arise at velocities km s-1 redward of NGC 4141. We obtain velocity-integrated Hi 21cm line flux densities of Jy km s-1 (NGC 4141), Jy km s-1 (Hi cloud), and Jy km s-1 (LEDA 2582852) from the spatial regions identified in the SoFiA search, after correcting for the shape of the GMRT primary beam. Using the Tully-Fisher distance estimate of Mpc (Tully et al., 2023) for NGC 4141, and assuming that the Hi cloud and LEDA 2582852 are at the same distance, these line flux densities yield Hi masses of (NGC 4141), (Hi cloud), and (LEDA 2582852). We note that the above uncertainties do not include systematic uncertainties in the flux density scale, which are typically % for the GMRT with the above calibration approach. Our measurement of the Hi mass of NGC 4141 is in reasonable agreement with that of Springob et al. (2005), who obtain a velocity-integrated Hi 21cm line flux density of Jy km s-1 [, from systematic uncertainties in the flux density scale] from an Effelsberg Telescope spectrum, after correcting for beam attenuation and pointing offsets, but not for possible self-absorption.
| NGC 4141 | Hi cloud | LEDA 2582852 | |
| Right Ascension (J2000) | 12h09m47.32s | 12h09m25.86s | 12h08m58.73s |
| Declination (J2000) | +58d50’5707 | +58d50’257 | +58d47’3166 |
| Redshift | 0.0063 | 0.0066 | 0.0064 |
| Tully-Fisher distance (Mpc) | – | – | |
| Separation (kpc) | – | 27 | 68 |
| (Jy km s-1) | |||
| SFR () | – | ||
| sSFR (yr-1) | – | ||
| 1.4 | – | 9.4 | |
| (Gyr) | 3.9 | – | 22.4 |
NGC 4141 has a stellar mass of and a star-formation rate (SFR) of year-1 (for a Chabrier initial mass function; Chabrier, 2003; Kewley et al., 2005; Leroy et al., 2019), i.e. a specific star-formation rate (sSFR) of yr-1. This implies an Hi-to-stellar mass ratio of and an Hi depletion timescale of Gyr. These are both within the spread of the corresponding values for galaxies of the xGASS sample with similar stellar masses (e.g. Saintonge et al., 2017; Catinella et al., 2018). Conversely, the companion galaxy, LEDA 2582852 has an Hi-to-stellar mass ratio of and an Hi depletion timescale of Gyr; these indicate that LEDA 2582852 is a relatively gas-rich galaxy. The Hi and stellar properties of NGC 4141, the Hi clump, and LEDA 2582852 are summarized in Table 2.
The four panels of Figure 3 show the first three velocity moments of the Hi 21cm image of NGC 4141, at an angular resolution of , produced by SoFiA searches using the smooth-and-clip algorithm (Serra et al., 2012, 2015). Fig. 3[A] shows the moment-0 image (in contours and colourscale), while Fig. 3[B] shows the moment-0 image (in contours) overlaid on the DECaLS g-band image (in colourscale). In these figures, the highest Hi column densities are seen to arise from the central region of the galaxy, and the two prominent spiral arms. However, high Hi column densities are seen to also arise in the south-west of the galaxy, where not much stellar emission is seen in the DECaLS image. The Hi spatial distribution does not appear symmetric, with extended Hi wings clearly visible to the north, the east, and the north-east of the galaxy. An Hi hole, with column densities of cm-2 can also be seen, just above the lower spiral arm and north-east of the galaxy centre. All of these signs of disturbances in the Hi spatial distribution suggest that the galaxy may have undergone a recent interaction. We note that the Hi 21cm moment-0 image of LEDA 2582852 of Figure 1[C] shows extensions to the east and north, also suggesting a recent interaction. Conversely, the Hi 21cm velocity field of NGC 4141, shown in Figure 3[C], appears broadly consistent with that expected from a rotating disk galaxy, albeit with some evidence of a turnover in velocity on the south-western side of the galaxy. Finally, the Hi 21cm velocity dispersion image of Fig. 3[D] shows that the highest velocity dispersions, of km s-1, are seen around the northern Hi extension, around the galaxy centre (in the direction of FRB 20250316A), and around the X-ray transient of Sun et al. (2025a, slightly south of FRB 20250316A). These velocity dispersions are significantly higher than those typically seen in disk galaxies (except for the central regions, e.g. Walter et al., 2008; de Blok et al., 2024). We note that the high observed velocity dispersions could arise due to beam-smearing, i.e. the crowding of iso-velocity contours within our angular resolution. Indeed, it is apparent from Figure 3[C] that the regions that show high velocity dispersions also show crowding of the iso-velocity contours. However, as discussed in detail by de Blok et al. (2024) for the MHONGOOSE survey (for spatial resolutions similar to ours), such crowding of iso-velocity contours is not expected to be significant away from the central regions of galaxies. The observed high velocity dispersions and the crowding of iso-velocity contours in the outer regions of NGC 4141 suggest that the galaxy is likely to have undergone a recent interaction.
Figure 4[A] shows the GMRT velocity-integrated Hi 21cm moment-0 image (contours) at a resolution of 30, overlaid on the DECaLS g-band image (colourscale; Dey et al., 2019). The top right panel, [B], shows a zoom-in on the region indicated by the dashed maroon square in [A]. While high column density gas can be seen north of, and close to, the location of FRB 20250316A, no Hi 21cm emission is detected at the FRB location (see panel [B]). We also extracted a spectrum at the location of FRB 20250316A from the 30-resolution cube, at a velocity resolution of km s-1; the spectrum (not shown here) shows no evidence for Hi 21cm emission. Our non-detection of Hi 21cm emission at the FRB location yields a upper limit on the local Hi column density of cm-2, at a spatial resolution of pc. We note that Hi 21cm emission is also not detected from the region immediately south of FRB 20250316A, where Chime/FRB Collaboration (2025b) identify a star-forming clump; it appears that the bulk of the Hi in this region has been either ionized or converted to molecular hydrogen.
| Line | Flux |
| erg s-1 cm-2 | |
| [Oii]3727Å | |
| H | |
| [Oiii]4363Å | |
| H | |
| [Oiii]4959Å | |
| [Oiii]5007Å | |
| H | |
| [Nii]6584Å |
Remarkably, Figures 4[A] and [B] show that high Hi column density gas is also detected in the south-west of NGC 4141, in the outer disk, which is quite faint in the DECaLS g-band image (see also Fig. 3[A] and [B]). Such high Hi column densities, cm-2, are typically found in the central, star-forming regions of galaxies, and not in the outskirts. We note that the high Hi column density gas in the south-western disk has a rough “V”-shape, with the western arm of the “V” terminating close to the bright star-forming knot visible in the DECaLS g-band image (and the X-ray transient of Sun et al. 2025a).
Figure 4[C] shows a Sloan Digital Sky Survey (SDSS) extended Baryon Oscillation Spectroscopic Survey (eBOSS; Dawson et al., 2016; Ahumada et al., 2020) spectrum from the location indicated by the magenta plus symbol in Figure 4[B], close to the highest Hi column density in the south-western region of the galaxy. The eBOSS spectrum shows strong [Oiii]5007Å emission, and weaker H and [Oii]3727Å emission. We emphasize that the central regions of NGC 4141 and the region around FRB 20250316A show stronger H and [Oii]3727Å lines, and weaker [Oiii]5007Å emission (e.g. Chime/FRB Collaboration, 2025b), typical of a normal star-forming galaxy; it is only the south-western region that shows the strong [Oiii]5007Å emission. Table 3 lists the observed flux densities of a number of the lines detected in the eBOSS spectrum.
To determine the SFR, the metallicity, etc., from the optical lines, one needs to first correct the line flux densities for dust extinction (see Chime/FRB Collaboration, 2025b, for a similar analysis for the FRB environment). We obtain a Galactic E(B-V) value of for this direction from the IRSA dust extinction database222http://irsa.ipac.caltech.edu/applications/DUST (Schlafly and Finkbeiner, 2011), corresponding to a low visual extinction, for . We hence ignore the Galactic extinction, effectively simply including it in the extinction internal to NGC 4141. We assume an intrinsic ratio of H to H line luminosities of 2.87 (e.g. Osterbrock and Ferland, 2006), and use the measured line flux densities to determine the visual extinction , assuming the G23 extinction model of Gordon et al. (2023b) (see also Gordon et al., 2009, 2021; Fitzpatrick et al., 2019; Decleir et al., 2022) and . This yields .
Applying the above value and the G23 extinction model to our measured H line flux density, we obtain a dust-corrected H line luminosity of erg s-1. Using the calibration of Kennicutt and Evans (2012), this yields an SFR of yr-1 at the location of the eBOSS spectrum, kpc from the centre of the galaxy, for a Chabrier initial mass function. The eBOSS fibres are circular, with diameter , implying an SFR surface density of yr-1 kpc-2 for the region covered by the eBOSS fibre. This is an uncharacteristically high SFR surface density for the outskirts of a disk galaxy, and suggests that the recent burst of star-formation is likely to have been triggered by a merger or accretion event.
Figure 5[A] shows the Galaxy Evolution Explorer (GALEX; Martin et al., 2005) near-ultraviolet (NUV) image of NGC 4141, in colourscale, with the velocity-integrated Hi 21cm emission from the -resolution cube overlaid in contours. Panel [B] of the figure shows a zoom-in on the dashed maroon square shown in [A], without the Hi 21cm contours. While NUV emission is indeed detected around the location of the eBOSS spectrum (indicated by the arrow), this region does not stand out in any way. The measured NUV flux at the above location is erg s-1 cm-2 Å-1, after convolving the image to an angular resolution of , where the errors are from Poisson statistics. For the G23 extinction model and , this yields an SFR of yr-1. This is significantly lower than the SFR of yr-1 from the same location from the H line. The apparent discrepancy can be explained by the fact that the mean age of the stellar population that contributes to the H line emission is Myr, while that contributing to the NUV continuum emission is Myr (e.g. Kennicutt and Evans, 2012). Thus, a very recent episode of star-formation, on timescales Myr would yield a higher SFR from the H line than the estimate from the NUV continuum. The lower SFR estimated from the GALEX NUV image of NGC 4141 is thus consistent with a burst of star-formation, triggered by a recent merger or accretion event. In passing, we note that the angular resolution of the GALEX NUV image is a factor of larger than the diameter of the eBOSS fibres. The SFR inferred from the NUV image is thus averaged over a larger region than that traced by the H emission.
We next used the “direct-Te” method to infer the metallicity at the location of the eBOSS spectrum (e.g. Izotov et al., 2006; Jiang et al., 2019), using the extinction-corrected flux densities in the [Oiii]4363Å, [Oiii]4959Å, [Oiii]5007Å, and [Oii]3727Å lines. This yields . The metallicity at the galaxy centre is (Kewley et al., 2005). This would imply a metallicity gradient of dex kpc-1 towards the south-west. Typical metallicity gradients in disk galaxies are dex kpc-1 (e.g. Zaritsky et al., 1994). A similar gradient, dex kpc-1, is obtained in NGC 4141 from the slightly sub-solar metallicity measured by Chime/FRB Collaboration (2025b) in the FRB environment. Therefore, the low metallicity at the location of the eBOSS spectrum instead suggests a recent addition of metal-poor gas, via either a merger or accretion from the circumgalactic medium. This fresh gas is likely to have resulted in the increased Hi column density seen in the south-western outskirts of the galaxy in Figs. 4[A] and [B], and to have triggered the recent star-formation activity, leading to the high inferred SFR surface density.
Leroy et al. (2019) find that NGC 4141 lies dex above the local star-forming main sequence. Our estimate of the 1.38 GHz flux density ( mJy) allows us to obtain an independent radio-based estimate of the SFR of NGC 4141. Using the Tully-Fisher distance estimate of Mpc (Tully et al., 2023) for the galaxy, we obtain a radio-based SFR of year-1. This is lower than the SFR estimate of year-1 from the dust-corrected H line luminosity (Kewley et al., 2005). As noted by Chime/FRB Collaboration (2025b), this difference may arise due to a recent increase in the star-formation activity, as the radio SFR estimate is an average over 100 Myr, while the H-based estimate is an average over the last 10 Myr. The increased recent SFR may account for the detections of two Type-II SNe, an optical transient, an X-ray transient, and FRB 20250316A, all within the last two decades, as well as for the fact that the galaxy lies above the local main sequence.
Overall, our GMRT Hi 21cm images find evidence for disturbances in the Hi distribution of NGC 4141 and the companion, LEDA 2582852, a nearby Hi cloud to the south-west without any discernible star-formation and at the velocities of the south-western side of the galaxy disk, and high Hi column densities in the south-western outskirts of the galaxy disk of NGC 4141. The SDSS spectrum of the south-western region of the galaxy yields a high SFR surface density and a low metallicity, while the recent ( Myr) SFR of NGC 4141 is higher than the radio-based SFR averaged over 100 Myrs. Further, the SFR inferred from the GALEX NUV image at the location of the eBOSS spectrum is significantly lower than that inferred from the H line. All of these indicate that NGC 4141 has recently acquired fresh gas, via either accretion or a merger, which compressed the Hi in the south-western part of the galaxy, triggering the starburst. This merger/accretion event may have also triggered additional star-formation activity in the disk of NGC 4141, resulting in the increased recent SFR (from the comparison between the radio and the H SFRs of the galaxy) and possibly in the formation of the stellar progenitor of FRB 20250316A, and the progenitors of the other transients. Earlier Hi 21cm mapping studies of FRB host galaxies have also yielded evidence for recent merger activity in a number of systems (e.g Kaur et al., 2022; Glowacki et al., 2024; Roxburgh et al., 2025), although we note that Roxburgh et al. (2025) also found cases where the FRB host galaxy showed no clear evidence for a merger. It is possible that the merger (merger or accretion event, in the case of FRB 20250316A) triggered a burst of star-formation activity that resulted in the formation of the massive stellar progenitors of the ensuing fast radio bursts. We emphasize, however, that the possibility that the progenitor of FRB 20250316A was formed independent of the merger/accretion event cannot be ruled out.
Finally, our non-detection of continuum emission at the location of FRB 20250316A places a upper limit of 24 Jy on the 1.38 GHz flux density of any associated persistent radio source (PRS). This implies the upper limit of erg s-1 on the rest-frame 1.38 GHz radio luminosity of such a putative PRS. This is one of the strongest constraints on the radio luminosity of a PRS associated with a non-repeating FRB (e.g. Law et al., 2024b; Chime/FRB Collaboration, 2025b; An et al., 2025), and approximately four orders of magnitude lower than the radio luminosities of the four repeating FRBs that show a PRS (e.g. Chatterjee et al., 2017; Niu et al., 2022; Bruni et al., 2025; Zhang et al., 2025). For example (for our assumed Tully-Fisher distance of 33.2 Mpc to NGC 4141), An et al. (2025) obtained erg s-1, while Chime/FRB Collaboration (2025b) obtained erg s-1. As noted by An et al. (2025), the stringent constraints on the flux density of a putative PRS associated with FRB 20250316A suggest, for pulsar wind nebula models (e.g. Dai et al., 2017), an evolved progenitor in a low-density, weakly magnetized environment, rather than a young magnetar in a dense environment.
IV Summary
We have used the GMRT Band-5 receivers to map Hi 21cm emission in and around NGC 4141, the host of FRB 20250316A, at , at angular resolutions of , i.e. spatial resolutions of kpc. The low redshift of NGC 4141 implies that this is the finest spatial resolution at which Hi 21cm emission has ever been mapped from an FRB host galaxy. The GMRT Hi 21cm images find evidence for a companion galaxy, LEDA 2582852, to the south-west of NGC 4141, a nearby Hi cloud also towards the south-west, high Hi column densities in the south-western outskirts of the galaxy disk, disturbances in the Hi distribution in both NGC 4141 and LEDA 2582852, and high velocity dispersions, km s-1, away from the central regions of NGC 4141. All of these suggest a recent merger or accretion event that compressed the Hi in the south-western disk, and resulted in a burst of star-formation activity in the galaxy. Consistent with this, an SDSS spectrum in the south-west yields evidence of a low metallicity, suggesting recent addition of metal-poor gas, via either accretion or a merger. The estimate of the recent SFR in NGC 4141, from the H line, is also higher than our estimate of the radio-based SFR, averaged over 100-Myr timescales, while the SDSS spectrum finds a high SFR surface density in the south-western disk. Finally, the SFR estimate from the H line in the south-western disk is significantly larger than the SFR estimate from the NUV continuum, indicating a very recent burst of star-formation, within the last Myr. We conclude that a merger or accretion event within the last Myr is likely to have compressed the Hi in the south-west of the galaxy and triggered the starburst activity. It is unclear whether the merger or accretion event was with the Hi cloud identified kpc away from NGC 4141 or whether this cloud was ejected from NGC 4141 by the merger/accretion event. Our highest-resolution Hi 21cm images, at a spatial resolution of pc at , find no evidence of Hi at the location of FRB 20250316A or in the neighbouring regions that show star-formation in recent KCWI images. While the location of FRB 20250316A is away from the location of the south-western starburst, the merger/accretion event is likely to have triggered star-formation activity across the galaxy (resulting in the higher galaxy-wide SFR estimate from the H line compared to the radio-based SFR estimate), possibly causing the formation of the massive stellar progenitor of FRB 20250316A. We obtain a upper limit of erg s-1 on the rest-frame 1.38 GHz luminosity of a persistent radio source associated with FRB 20250316A, one of the strongest current constraints on the radio luminosity of a PRS associated with a non-repeating fast radio burst.